# Solar flares in the Solar Orbiter Era: Short exposure EUI/FSI observations of STIX flares

### v1

@@ Abstract
Aims. This paper aims to demonstrate the importance of short-exposure extreme ultraviolet (EUV) observations of solar flares in the study of particle acceleration, heating and energy partition in flares. This work highlights the observations now available from the Extreme Ultraviolet Imager (EUI) instrument suite onboard Solar Orbiter while operating in short exposure mode.
Methods. A selection of noteworthy flares observed simultaneously by the Spectrometer Telescope for Imaging X-rays (STIX) and the Full Sun Imager of EUI (EUI/FSI) are detailed. New insights are highlighted and potential avenues of investigation are demonstrated, including forward modelling the atmospheric response to a non-thermal beam of electrons using the RADYN 1D hydrodynamic code, in order to compare the predicted and observed EUV emission.
Results. The examples given in this work demonstrate that short exposure EUI/FSI observations are providing important diagnostics during flares. A dataset of more than 9000 flares observed by STIX (from November 2022 until December 2023) with at least one short exposure EUI/FSI 174 Å image is currently available. The observations reveal that the brightest parts of short-exposure observations consist of substructure in flaring ribbons which spatially overlap with the hard X-ray emission observed by STIX in the majority of cases. We show that these observations provide an opportunity to further constrain the electron energy flux required for flare modelling, among other potential applications.
@@ 1 Introduction
During solar flares, the emission at ultraviolet (UV) and extreme ultraviolet (EUV) wavelengths can become locally enhanced by several orders of magnitude on a timescale of seconds to minutes. Even moderately sized C-class flares typically saturate current EUV and UV imagers as well as soft X-ray imagers. This means that spatial information in the saturated pixels is lost. Further, depending on the type of detectors used (CMOS vs. CCD), the total flux may or may not be conserved and neighbouring pixels might be affected (so called “bleeding”). Saturation during enhanced solar flare emission in such instruments is due to the fact that they are not exclusively designed to study flares. It is challenging to design an instrument with a sufficiently large dynamic range that can observe both the faintest solar emission and the largest flares for a given exposure time. The saturation effect leads to challenges in many flare studies, particularly for relating coronal and chromospheric UV/EUV emission to observations from hard X-ray emission during the impulsive phase of flares.
@@
One way to address the saturation issue of EUV imagers is to shorten the exposure time used. Some EUV/UV imagers, such as Atmospheric Imaging Assembly (AIA) onboard the Solar Dynamics Observatory, use automatic exposure control, in which flight software determines the exposure time based on count flux (Lemen et al., 2012). However, with variations in this exposure time over the course of a flare, it can become a challenge to study given that the response of the detectors may be non-linear in exposure time which can introduce artificial variations linked to the exposure time changes. Regardless, despite the variable exposure time setting, the exposure time is often too long and AIA observations of flares are frequently still saturated. Efforts have been made to overcome saturation issues through a reconstruction approach which utilises the knowledge of the AIA diffraction pattern (Raftery et al., 2011; Krucker et al., 2011; Schwartz et al., 2014; Torre et al., 2015; Guastavino et al., 2019; Krucker et al., 2021). However, there are still challenges with total flux conservation and dealing with pixel bleeding. Kazachenko et al. (2017) also tackled this issue by linearly interpolating the flux prior to and post saturation in order to infer spatial information about flare ribbons. An approach of this kind, is, however, not suitable for photometric analysis.
@@
In the era of multi-messenger observations, the availability of EUV imagers on the far side of the Sun from Earth has become increasingly important for event studies since we cannot rely on Earth-based assets like AIA. For example, the Extreme Ultraviolet Imager (EUI) (Rochus et al., 2020; Berghmans et al., 2023) onboard Solar Orbiter (Müller et al., 2020) provides EUV images for observations taken by the Solar Orbiter instrument suite as the spacecraft sweeps out an elliptical orbit spending half of the mission duration on the far side of the Sun from Earth. EUI consists of three telescopes, the Full Sun Imager (EUI/FSI) and two High Resolution Imagers, which observe in Lyman , 174 Å and 304 Å passbands. The focus of this work is on EUI/FSI observations.
@@
To combat the aforementioned saturation issues, EUI/FSI has been regularly taking short exposure observations alongside normal exposure frames in both the 174 Å and 304 Å channels since the end of 2022. This paper demonstrates the value of short exposure EUV observations in flare studies and details the invaluable new datasets provided by EUI/FSI as well as recent observations with EUI’s High Resolution Imager, EUI/HRIEUV . In this work, a subset of interesting flares co-observed by the Spectrometer Telescope for Imaging X-rays (STIX) are presented and new insights enabled by this observation mode are discussed.
@@ 2.1 EUI/FSI and the short-exposure observing mode
EUI/FSI is an EUV imager observing the full solar disk in 174 Å and 304 Å EUV passbands. It continually provides synoptic observations with a variable cadence depending on telemetry allocations and observing mode. During FSI synoptic imaging, the two bandpasses are alternated by rotating a filter wheel (Rochus et al., 2020). Since 09-Nov-2022 EUI/FSI has been regularly taking short exposure observations alongside normal exposure (see Fig 1). The nominal exposure time for an FSI image in either bandpass is 10 s. Reading out a full sensor FSI frame (3072x3072 pixels) takes about 2.6 s seconds. Immediately before the regular exposure, two more images are taken during a very brief exposure time and read-out. The first of these is a dummy image that is disregarded upon arrival in the instrument computer. Taking a dummy image resets the sensor to a stable, reproducible status, independent from the previous history. The second extra image is the short-exposure image which is intended to record intensity values where the (third) regular image is saturated. We therefore choose a significantly shorter exposure time (1/50th, i.e. 0.2 s). Regular exposure FSI images contain intensity values in the range 0–32767 DN. Unsaturated pixels (i.e. with values DN) are not needed in the short-exposures as they are recorded in the regular image. Leaving a factor 2 margin, we therefore limit the intensity range in the short exposures to values above DN = 327 DN.
The cadence of the EUI/FSI synoptics varies as a function of telemetry and can be between several minutes and half an hour. The observing settings described here are merely examples as the instrument is highly configurable and can be adjusted for different use-cases. The main constraint of deep space missions like Solar Orbiter is limited telemetry, however, as discussed, for short exposure images, only the brightest pixels, with pixel counts in the upper half of the dynamic range of regular images, are additionally downloaded. The other, dimmer and noisier pixels all get the same minimal value in the short exposure image (327 DN), which facilitates compression of the short exposure images greatly. Typically, the short exposure data volume corresponds to of the data volume of normal exposures for EUI/FSI images and as little as for EUI/HRIEUV frames.
@@ 2.2 The STIX instrument
STIX is a hard X-ray (HXR) imaging spectrometer measuring emission in the energy range 4-150 keV, from thermal and non-thermal flare accelerated electrons (Krucker et al., 2020). STIX observes the full solar disk at a minimum cadence of 0.5 s and operates continuously, unlike the other remote-sensing instruments on Solar Orbiter which primarily operate during planned operational windows. STIX has been taking science data since early 2021, and to date (July 2024), STIX has observed more than 50,000 flares. It is an indirect Fourier imager meaning that images are reconstructed from Fourier components sampled by the instrument (STIX samples 30 Fourier components or visibilities). There are certain constraints that arise from this fact including a limited instrument dynamic range. This means that STIX is capable of only resolving the brightest X-ray emission. This limitation depends on counting statistics and relative brightness. For more information about the STIX imaging concept, we refer the reader to Massa et al. (2023).
STIX provides us with critical diagnostics of solar flare energy release and heating through thermal and non-thermal HXR emission. Complementary EUV flare observations provide us with insight into heated flare plasma with both high resolution and dynamic range. In particular, STIX observations give insight into the released flare energy which is transferred to high-energy electrons whereas EUI/FSI 304 Å and 174 Å give insight into the chromospheric and coronal response to flare energy release, deposition and heating. EUI provides snapshots of flares and flaring active regions that are critical for STIX observations especially at times when Solar Orbiter is on the far side of the Sun from Earth and therefore Earth based context observations are not available. Additionally, short exposure EUV observations help to verify the sources reconstructed with STIX imaging algorithms such as Clean (Hurford et al., 2002), MEM_GE (Massa et al., 2020) and forward-fit PSO (Volpara et al., 2022), the presence of which may otherwise be inconclusive due to dynamic range limitations. Furthermore, the highly resolved flare kernels obtained from EUV flare observations provide important constraints on flare energy deposition rates.
@@ 3 Available data and example case studies
A common flarelist has been created of STIX flares with short exposure EUI/FSI observations. This list was compiled from the operational STIX flarelist available at the STIX data center111https://datacenter.stix.i4ds.net/ and is also accessible via the API of the python library stixdcpy222https://github.com/i4Ds/stixdcpy (Xiao et al., 2023). Flares with ¿ 1000 peak counts in the 10-15 keV quicklook channel were considered. The available Level 2 (L2) EUI/FSI 174 Å short exposure files within the flare time range were obtained through the SIDC website333https://www.sidc.be/EUI/data/, these data are also available from https://soar.esac.esa.int/soar/. From this, a combined list of STIX-EUI/FSI short exposure observations was created. This list considers the time range 09-Nov-2022 - 31-Nov-2023 (the date from which short exposure observations began until the most recent publicly available L2 EUI files at the time of analysis). We find that % of STIX flares in this time range have at least one short exposure frame which amounts to 9481 co-observed flares; the code used to create this list is publicly available444https://github.com/hayesla/eui_gi_visit. From this flarelist a subset of events have been further analysed and three events which demonstrate the advantages of this unique data set are presented in the sections which follow.
@@ 3.1 STX2023-07-16T04:32: Rapid pulsations
STX2023-07-16T04:32555The STX flare identifier used throughout this work labels each flare by the date and time at which the flare counts reached a maximum in the STIX 4-10 keV quicklook channel. is one of the largest flares observed by STIX to date. It was estimated to be of GOES class X9 using the method proposed by Xiao et al. (2023). At the time Solar Orbiter was at to the Sun-Earth line and as the flare was located at , in heliographic stonyhurst coordinates, the flare was not observable from Earth. The flare displayed rapid variation on timescales of 10 s in the non-thermal HXR emission (see Fig. 2, panel a), making it an interesting case to study quasi-periodic pulsations and rapid variation in flaring emission (e.g. Collier et al., 2023, 2024). EUI/FSI observed this flare in both short and normal exposure mode after the non-thermal peak time but before the thermal peak. The time of the short exposure is denoted by the vertical dashed line in Fig. 2 a). In the normal exposure frame (panel b), the flare site is saturated, however the short exposure (panel c) distinctly shows the flaring ribbons and the fine structure within i.e. the short exposure frame highlights the distribution of brightness along the length of the northern ribbon as well as the spatially compact bright source in the southern flare ribbon. Due to a known non-linearity in the instrumental response with exposure time, a correction factor is required to match the fluxes observed by the normal and short exposures (E. Kraaikamp 2024, priv. communication). The cause of the non-linearity is currently being investigated. In this case the flux correction factor used was 1.33. This factor was obtained by matching the flux in non-saturated pixels in the regular exposures with the corresponding pixels in the short exposures for a selection of flares. It is therefore effectively a shortening of the exposure time by this factor. Given the current level of understanding of this discrepancy, it is recommended that this analysis should be performed on a case-by-case basis. Panel d) shows a composite map of the short and normal exposure frames. It has been rotated such that Solar north is “up” for comparison with the STIX contours. Since the counts in the short exposure maps are spatially sparse, performing accurate rotations can be tricky as it typically involves interpolation between pixels. In this work, the rotation functionality provided by sunpy (The SunPy Community et al., 2020), sunpy.map.GenericMap.rotate, is used. In particular, the “scikit-image” method was used as the default rotation method (“scipy”) uses spline interpolation which involves a pre-filtering step and this step causes discontinuities to have ripple-like effects. This effect becomes particularly prominent with short exposure maps as there are sharp spatial discontinuities in flux. In panel d) the non-thermal (36-76 keV) and thermal (6-15 keV) HXR emission is shown for the time range 04:30:40-04:31:20 UT - shaded in blue in panel a) (all times given in this work are at the observing telescope). A shift of (-10, 35)” has been applied to align the STIX maps to the short exposure frame. Interestingly, the separation of the HXR non-thermal sources corresponds remarkably well with that of the ribbons revealed by the short exposure EUI/FSI frame. The thermal loop appears to sit on the northern ribbon, however, this is likely due to projection effects - the flare was located at approximately from the central meridian while Solar Orbiter was at . The 45-76 keV light curve of this flare shows rapid temporal variation. This temporal variability is a signature of fundamental energy release and particle acceleration processes ongoing in flares and therefore is a crucial diagnostic required to reach a unified flare model. The spatial evolution of the atmospheric response to this electron beam variability is currently unresolved temporally, given the low cadence EUI/FSI observations while in synoptic mode. While these synoptic observations are vitally important for understanding the surrounding active region structure, the need for high time cadence short exposure observations of flaring ribbons in EUV is apparent.
3.2 STX2022-11-13T06:18: A “standard” flare

### v2

# Solar flares in the Solar Orbiter era: Short-exposure EUI/FSI observations of STIX flares

@@ Abstract
Aims. This paper aims to demonstrate the importance of short-exposure extreme ultraviolet (EUV) observations of solar flares in the study of particle acceleration, heating, and energy partition in flares. This work highlights the observations now available from the Extreme Ultraviolet Imager (EUI) instrument suite on board Solar Orbiter while operating in short-exposure mode.
Methods. A selection of noteworthy flares observed simultaneously by the Spectrometer Telescope for Imaging X-rays (STIX) and the Full Sun Imager of EUI (EUI/FSI) are detailed. New insights are highlighted and potential avenues of investigation are demonstrated, including forward-modelling of the atmospheric response to a non-thermal beam of electrons using the RADYN 1D hydrodynamic code, in order to compare the predicted and observed EUV emission.
Results. The examples given in this work demonstrate that short-exposure EUI/FSI observations are providing important diagnostics during flares. A dataset of more than 9000 flares observed by STIX (from November 2022 until December 2023) with at least one short-exposure EUI/FSI 174 Å image is currently available. The observations reveal that the brightest parts of short-exposure observations consist of substructure in flaring ribbons that spatially overlap with the hard X-ray emission observed by STIX in the majority of cases. We show that these observations provide an opportunity to further constrain the electron energy flux required for flare modelling, among other potential applications.
@@ 1 Introduction
During solar flares, the emission at ultraviolet (UV) and extreme ultraviolet (EUV) wavelengths can become locally enhanced by several orders of magnitude on a timescale of seconds to minutes. Even moderately sized C-class flares typically saturate current EUV and UV imagers as well as soft X-ray imagers. This means that spatial information in the saturated pixels is lost. Furthermore, depending on the type of detectors used, the total flux may or may not be conserved and neighbouring pixels might be affected (so called ‘bleeding’). Saturation during enhanced solar flare emission in such instruments is due to the fact that they are not exclusively designed to study flares. It is challenging to design an instrument with a dynamic range that is sufficiently large to observe both the faintest solar emission and the largest flares for a given exposure time. The saturation effect leads to challenges in many flare studies, particularly for relating coronal and chromospheric UV/EUV emission to observations from hard X-ray emission during the impulsive phase of flares.
@@
One way to address the saturation issue of EUV imagers is to shorten the exposure time used. Some EUV/UV imagers, such as the Atmospheric Imaging Assembly (AIA) on board the Solar Dynamics Observatory, use automatic exposure control, in which flight software determines the exposure time based on the count flux (Lemen et al., 2012). However, with variations in this exposure time over the course of a flare, it can become a challenge to study given that the response of the detectors may be non-linear in exposure time, which can introduce artificial variations linked to the exposure time changes. Regardless, despite the variable exposure time setting, the exposure time is often too long and AIA observations of flares are frequently still saturated. Efforts have been made to overcome saturation issues through a reconstruction approach that utilises knowledge of the AIA diffraction pattern (Raftery et al., 2011; Krucker et al., 2011; Schwartz et al., 2014; Torre et al., 2015; Guastavino et al., 2019; Krucker et al., 2021). However, there are still challenges with total flux conservation and dealing with pixel bleeding. Kazachenko et al. (2017) also tackle this issue by linearly interpolating the flux pre- and post-saturation in order to infer spatial information about flare ribbons. An approach of this kind is, however, not suitable for photometric analysis.
@@
In the era of multi-messenger observations, the availability of EUV imagers on the far side of the Sun from Earth has become increasingly important for event studies, since we cannot rely on Earth-based assets like AIA. For example, the Extreme Ultraviolet Imager (EUI) (Rochus et al., 2020; Berghmans et al., 2023) on board Solar Orbiter (Müller et al., 2020) provides EUV images for observations taken by the Solar Orbiter instrument suite as the spacecraft sweeps out an elliptical orbit, spending half of the mission duration on the far side of the Sun from Earth. EUI consists of three telescopes, the Full Sun Imager (EUI/FSI) and two High Resolution Imagers, which observe in Lyman , 174 Å , and 304 Å passbands. The focus of this work is on EUI/FSI observations.
@@
To combat the aforementioned saturation issues, EUI/FSI has been regularly taking short-exposure observations alongside normal-exposure frames in both the 174 Å and 304 Å channels since the end of 2022. This paper demonstrates the value of short-exposure EUV observations in flare studies and details the invaluable new datasets provided by EUI/FSI as well as recent observations with EUI’s High Resolution Imager, EUI/HRIEUV . In this work, we present a subset of interesting flares co-observed by the Spectrometer Telescope for Imaging X-rays (STIX) and discuss new insights enabled by this observation mode.
@@ 2.1 EUI/FSI and the short-exposure observing mode
EUI/FSI is an EUV imager observing the full solar disc in the 174 Å and 304 Å EUV passbands. It continually provides synoptic observations with a variable cadence depending on telemetry allocations and observing mode. During FSI synoptic imaging, the two bandpasses are alternated by rotating a filter wheel (Rochus et al., 2020). Since November 9 2022, EUI/FSI has been regularly taking short-exposure observations alongside normal-exposure ones (see Fig 1). The nominal exposure time for an FSI image in either bandpass is 10 s. Reading out a full sensor FSI frame (3072x3072 pixels) takes about 2.6 s seconds. Immediately before the regular exposure, two more images are taken during a very brief exposure time and read-out. The first of these is a dummy image that is disregarded upon arrival in the instrument computer. Taking a dummy image resets the sensor to a stable, reproducible status, independent from the previous history. The second extra image is the short-exposure image, which is intended to record intensity values where the (third) regular image is saturated. We therefore chose a significantly shorter exposure time (1/50th, i.e. 0.2 s). Regular exposure FSI images contain intensity values in the range of 0–32767 DN. Unsaturated pixels (i.e. with values DN) are not needed in the short-exposures as they are recorded in the regular image. Leaving a factor-two margin, we therefore limited the intensity range in the short exposures to values above DN = 327 DN.
The cadence of the EUI/FSI synoptics varies as a function of telemetry and can be between several minutes and half an hour. The observing settings described here are merely examples as the instrument is highly configurable and can be adjusted for different use-cases. The main constraint of deep space missions like Solar Orbiter is limited telemetry; however, as was discussed, for short-exposure images, only the brightest pixels, with pixel counts in the upper half of the dynamic range of regular images, are additionally downloaded. The other, dimmer and noisier pixels all get the same minimal value in the short-exposure image (327 DN), which facilitates compression of the short-exposure images greatly. Typically, the short-exposure data volume corresponds to of the data volume of normal exposures for EUI/FSI images and as little as for EUI/HRIEUV frames.
@@ 2.2 The STIX instrument
STIX is a hard X-ray (HXR) imaging spectrometer measuring emission in the energy range 4-150 keV, from thermal and non-thermal flare accelerated electrons (Krucker et al., 2020). STIX observes the full solar disc at a minimum cadence of 0.5 s and operates continuously, unlike the other remote-sensing instruments on Solar Orbiter, which primarily operate during planned operational windows. STIX has been taking science data since early 2021, and to date (July 2024), STIX has observed more than 50,000 flares. It is an indirect Fourier imager, meaning that images are reconstructed from Fourier components sampled by the instrument (STIX samples 30 Fourier components or visibilities). There are certain constraints that arise from this fact, including a limited instrument dynamic range. This means that STIX is capable of only resolving the brightest X-ray emission. This limitation depends on counting statistics and relative brightness. For more information about the STIX imaging concept, we refer the reader to Massa et al. (2023).
STIX provides us with critical diagnostics of solar flare energy release and heating through thermal and non-thermal HXR emission. Complementary EUV flare observations provide us with an insight into heated flare plasma with both high resolution and dynamic range. In particular, STIX observations give an insight into the released flare energy that is transferred to high-energy electrons, whereas EUI/FSI 304 Å and 174 Å give an insight into the chromospheric and coronal response to flare energy release, deposition, and heating. EUI provides snapshots of flares and flaring active regions that are critical for STIX observations, especially at times when Solar Orbiter is on the far side of the Sun from Earth, and therefore Earth-based context observations are not available. Additionally, short-exposure EUV observations help to verify the sources reconstructed with STIX imaging algorithms such as Clean (Hurford et al., 2002), MEM_GE (Massa et al., 2020), and forward-fit PSO (Volpara et al., 2022), the presence of which may otherwise be inconclusive due to dynamic range limitations. Furthermore, the highly resolved flare kernels obtained from EUV flare observations provide important constraints on flare energy deposition rates.
@@ 3 Available data and example case studies
A common flare list has been created of STIX flares with short-exposure EUI/FSI observations. This list was compiled from the operational STIX flare list available at the STIX data center111https://datacenter.stix.i4ds.net/ and is also accessible via the API of the Python library stixdcpy222https://github.com/i4Ds/stixdcpy (Xiao et al., 2023). Flares with ¿ 1000 peak counts in the 10-15 keV quicklook channel were considered. The available Level 2 (L2) EUI/FSI 174 Å short-exposure files within the flare time range were obtained through the SIDC website333https://www.sidc.be/EUI/data/, these data are also available from https://soar.esac.esa.int/soar/. From this, a combined list of STIX-EUI/FSI short-exposure observations was created. This list considers the time range from November 9 2022 to November 31 2023 (the date from which short-exposure observations began until the most recent publicly available L2 EUI files at the time of analysis). We find that % of STIX flares in this time range have at least one short-exposure frame, which amounts to 9481 co-observed flares; the code used to create this list is publicly available444https://github.com/hayesla/eui_gi_visit. From this flare list, a subset of events have been further analysed and three events that demonstrate the advantages of this unique dataset are presented in the sections that follow.
@@ 3.1 STX2023-07-16T04:32: Rapid pulsations
STX2023-07-16T04:32555The STX flare identifier used throughout this work labels each flare by the date and time at which the flare counts reached a maximum in the STIX 4-10 keV quicklook channel. is one of the largest flares observed by STIX to date. It was estimated to be of GOES class X9 using the method proposed by Xiao et al. (2023). At the time, Solar Orbiter was at to the Sun-Earth line and as the flare was located at , in heliographic stonyhurst coordinates, the flare was not observable from Earth. The flare displayed rapid variation on timescales of 10 s in the non-thermal HXR emission (see Fig. 2, panel a), making it an interesting case in which to study quasi-periodic pulsations and rapid variation in flaring emission (e.g. Collier et al., 2023, 2024). EUI/FSI observed this flare in both short- and normal-exposure mode after the non-thermal peak time but before the thermal peak. The time of the short exposure is denoted by the vertical dashed line in Fig. 2 a). In the normal exposure frame (panel b), the flare site is saturated; however, the short exposure (panel c) distinctly shows the flaring ribbons and the fine structure within; that is, the short exposure frame highlights the distribution of brightness along the length of the northern ribbon as well as the spatially compact bright source in the southern flare ribbon. Due to a known non-linearity in the instrumental response with exposure time, a correction factor is required to match the fluxes observed by the normal and short exposures (E. Kraaikamp 2024, priv. communication). The cause of the non-linearity is currently being investigated. In this case, the flux correction factor used was 1.33. This factor was obtained by matching the flux in non-saturated pixels in the regular exposures with the corresponding pixels in the short exposures for a selection of flares. It is therefore effectively a shortening of the exposure time by this factor. Given the current level of understanding of this discrepancy, it is recommended that this analysis be performed on a case-by-case basis. Panel d) shows a composite map of the short and normal exposure frames. It has been rotated such that Solar north is up for comparison with the STIX contours. Since the counts in the short-exposure maps are spatially sparse, performing accurate rotations can be tricky as this typically involves interpolation between pixels. In this work, the rotation functionality provided by sunpy (The SunPy Community et al., 2020), sunpy.map.GenericMap.rotate, has been used. In particular, the ‘scikit-image’ method was used as the default rotation method (‘scipy’) uses spline interpolation, which involves a pre-filtering step, and this step causes discontinuities to have ripple-like effects. This effect becomes particularly prominent with short-exposure maps as there are sharp spatial discontinuities in flux. In panel d) the non-thermal (36-76 keV) and thermal (6-15 keV) HXR emission is shown for the time range 04:30:40-04:31:20 UT – shaded in blue in panel a) (all times given in this work are at the observing telescope). A shift of (-10, 35)” has been applied to align the STIX maps with the short-exposure frame. Interestingly, the separation of the HXR non-thermal sources corresponds remarkably well to that of the ribbons revealed by the short-exposure EUI/FSI frame. The thermal loop appears to sit on the northern ribbon; however, this is likely due to projection effects – the flare was located at approximately from the central meridian, while Solar Orbiter was at . The 45-76 keV light curve of this flare shows rapid temporal variation. This temporal variability is a signature of fundamental energy release and particle acceleration processes ongoing in flares, and therefore is a crucial diagnostic required to reach a unified flare model. The spatial evolution of the atmospheric response to this electron beam variability is currently unresolved temporally, given the low-cadence EUI/FSI observations while in synoptic mode. While these synoptic observations are vitally important for understanding the surrounding active region structure, the need for high-time-cadence short-exposure observations of flaring ribbons in EUV is apparent.
3.2 STX2022-11-13T06:18: A ‘standard’ flare